OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, -5) with a slope of -1?

OpenStudy (anonymous):

(1,-5)_m=-1 y=mx+b (-5)=(-1)*(1)+b b=-4 y=-x-4 x+y=-4